OpenStudy (anonymous):

A=1/2 BH how would you solve for b and how did you get the answer?

OpenStudy (anonymous):

A = 1/2 BH means that the area of a triangle = 1/2 (base) * (height). To solve for B, you'd need both the area of the triangle, and the height of the triangle.

OpenStudy (anonymous):

But it's asking to re write the problem and to write it as though you were solving for B since we don't know what the numbers for the base or the height are.

OpenStudy (anonymous):

How u get two

OpenStudy (anonymous):

Sorry, it was wrong. Let me try to explain it again. A = 1/2BH Move height to the left which means dividing H on both sides. (A/H) = 1/2B Now multiply 2 on both sides to get rid of the 1/2 on the right. 2(A/H) = B

OpenStudy (anonymous):

Cool thank you